OpenStudy (anonymous):

Write an equation of a line in point-slope form with a slope of -3/4 and passing through the point (1,-5)

OpenStudy (anonymous):

I think point slope form is y=mx+b, but I couldnt be mistaken (kinda forgot the names of some things) so, y=mx+b m=-3/4 so y=-3/4x+b y=-5, x=1: -5=-3/4(1)+b b=-5+3/4 b=-17/4 y=-3/4x-17/4

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You are given the slope in this problem, so to find the equation of a line plug the slope, x, and y values into y-y1=m(x-x1)

OpenStudy (anonymous):

where slope = m

OpenStudy (anonymous):

and you will have the equation

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

@klerg10, you gave standard form point-slope form is y-y1=m(x-x1)

OpenStudy (anonymous):

@kierg10

OpenStudy (anonymous):

if you solve it down it gets to the same thing.

OpenStudy (anonymous):

yeah, still good to know the difference in standard form and slope form :)

OpenStudy (anonymous):

if you don't need any more help @cacique, please close the question

OpenStudy (anonymous):

y-y1=m(x-x1) y--5=-3/4(x-1) y+5=-3/4(x-1) y+5=-3/4x+3/4 y=-3/4x+3/4-5 y=-3/4x-17/4

OpenStudy (anonymous):

so it's the same thing XD

OpenStudy (anonymous):

you get to the same answer

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

IDK about you, but personally no teacher I have ever had has used the y-y1=m(x-x1) format......the first time I encountered it I was really confused.

OpenStudy (anonymous):

I have used the point slope formula many times That is why I have it down

OpenStudy (anonymous):

I know the standard form too but since the question asked for point-slope form that is what we used

OpenStudy (anonymous):

true